OpenStudy (anonymous):

Question: 53 of 63: Lisa invests $4,000 in two types of bonds: X, which offers a 10% return, and Y, which offers a 6% return. If the combined return on the investment is $300, how much did Lisa invest at the rate of 6%? Select one of the options below as your answer: A. $3,000 B. $2,500 C. $1,500 D. $500

OpenStudy (anonymous):

Lisa invests $4,000 in two types of bonds: X, which offers a 10% return, and Y, which offers a 6% return. If the combined return on the investment is $300, how much did Lisa invest at the rate of 6%? Let X be amount invested in the 10%. Amount invested in 6% = 4000 - X Thus, Return for 10% = 0.1X Thus, Return for 6% = 0.06(4000-X) Total Return = 300 = 0.1X + 0.06(4000-X) 300 = 0.1X + 240 - 0.06X 60 = 0.04X X = 60 / 0.04 = ????????

OpenStudy (anonymous):

1500?